OpenStudy (anonymous):

surface area of a cone with a slant height of 19 and a radius of 7

OpenStudy (jdoe0001):

 \(\textit{Total Surface of a Cone}=\boldsymbol{\pi r \cdot \textit{slant height}+\pi r}\)